OpenStudy (anonymous):

graphing the sq rt. functions: g(x)=sq rt x-1 (x=0,1,4,9) i understand to replace x with the numbers to come up with 4 relations, but my problem is after subtracting the 1 under the sq rt, my problem is removing the sq rt symbol, yes i know th sq the number, but how can you sq a (0,1,3)

OpenStudy (fifciol):

you can take only the approximate values

OpenStudy (anonymous):

so do i quit the problem and write "function cannot be defined"?

OpenStudy (anonymous):

for my answer

OpenStudy (anonymous):

my book has answers to this

OpenStudy (fifciol):

wouldn't you take for x =(1,5,10,..) ?

OpenStudy (anonymous):

hang on, i took a picture from my book, let me figure out how to get it to u

OpenStudy (fifciol):

ok

OpenStudy (anonymous):

unless u wanna give me ur # and i can just text to u, if not hang on

OpenStudy (anonymous):

look at the bottom, however it is the same problem

OpenStudy (anonymous):

√4-1=1.....how? √0-1=-1...how?

OpenStudy (fifciol):

oh, I thought it was \[\sqrt{x-1}\] but it's \[\sqrt x-1\] so for x=4 you have \[\sqrt4 -1=2-1=1\] for x =0: \[\sqrt0 -1=0-1=-1\] and so on

OpenStudy (anonymous):

ooooh, only first # is under the sq rt?

OpenStudy (fifciol):

yes

OpenStudy (anonymous):

i feeeeeel stttuuuuuuuupppppppiiiiiiidddddd

OpenStudy (fifciol):

I've been stupid so many times, but this motivates you to become better

OpenStudy (anonymous):

f(x)=x^4-x^2+1 x=-x when replacing is it x^4-(-x)+1 or x^4-x+1

OpenStudy (fifciol):

\[(-x)^4 -(-x)^2+1\]

OpenStudy (fifciol):

but you should have put it on new post

OpenStudy (anonymous):

y u got - in front of x^4?

OpenStudy (fifciol):

I replaced x with -x

OpenStudy (anonymous):

original dont have it

OpenStudy (anonymous):

no on first variable, maybe u made mistake

OpenStudy (anonymous):

f(x)=x^4-x^2+......see

OpenStudy (fifciol):

but (-x)^4 is simply x^4

OpenStudy (anonymous):

ooh shucks, sorry, im trippy mane!

OpenStudy (anonymous):

new post time but it picks up from here!